There are 20 questions in a multiple-choice test. Each question has five choices and one correct answer out of these five choices. A student did not study. He/she answers the questions at random What are the probabilities that he/she makes (i) more than 10-answers right? (ii) exatly eight right? And (iii) at most 7 right?
Answer by stanbon(48569)

There are 20 questions in a multiple-choice test. Each question has five choices and one correct answer out of these five choices. A student did not study. He/she answers the questions at random What are the probabilities that he/she makes
(i) P(more than 10-answers right) = 1-binomcdf(20,(1/5),10) = 0.0005634
(ii) P(exactly eight right) = binompdf(20,5^-1,8) = 0.0222...
(iii) P(at most 7 right) = binomcdf(20,5^-1,7) = 0.9679..
Cheers,
Stan H.